# MagicTailor: Component-Controllable Personalization in Text-to-Image Diffusion Models

### v1

@@ 4.5 Further Applications
Decoupled Generation. After learning from a concept-component pair, MagicTailor can also enable decoupled generation. As shown in Figure 8(a), MagicTailor can generate the target concept and component separately in various and even cross-domain contexts. This should be credited to its remarkable ability to accurately capture different-level visual semantics. Such an ability extends the flexibility of the possible combination between the concept and component.
@@
Controlling Multiple Components. In this paper, we focus on personalizing one concept and one component, because such a setting is enough to cover extensive scenarios in the real world, and can be further extended to reconfigure multiple components with an iterative procedure. However, as shown in Figure 8(b), our MagicTailor also exhibits the potential to handle one concept and multiple components simultaneously. These results reflect a prospective direction of exploring better control over diverse components for a single concept.
@@
Enhancing Other Generative Tools. In Figure 9, we provide some interesting method combinations to show that our MagicTailor can enhance other generative tools. The combined tools include ContorlNet (Zhang et al., 2023), CSGO (Xing et al., 2024), and InstantMesh (Xu et al., 2024). As we can see, MagicTailor can be seamlessly integrated into these tools, furnishing them with an additional ability to control the concept’s component in their pipelines. For instance, working with MagicTailor, InstantMesh can conveniently achieve fine-grained design of 3D mesh, demonstrating the practicability of MagicTailor in collaborative applications.
Figure 9: Enhancing other generative tools. MagicTailor can conveniently collaborate with a variety of generative tools that focus on other tasks, equipping them with an additional ability to control the concept’s component in their pipelines.
@@ 5 Conclusion
In this paper, we introduce the novel task of component-controllable personalization, which allows for precise customization of individual components within a personalized concept. We tackle two major challenges that make this task particularly difficult: semantic pollution, where unwanted visual elements disrupt the integrity of the concept, and semantic imbalance, which skews the learning process of visual semantics. To address these challenges, we present MagicTailor, an innovative framework featuring Dynamic Masked Degradation (DM-Deg) to mitigate unwanted visual semantics and Dual-Stream Balancing (DS-Bal) to ensure balanced learning of visual components. Our comprehensive experiments demonstrate that MagicTailor not only sets a new benchmark in this challenging task but also opens up exciting possibilities for a wide range of creative applications. Looking ahead, we envision extending our approach to other areas of image and video generation, exploring how multi-level visual semantics can be recognized, controlled, and manipulated to unlock even more sophisticated and imaginative generative capabilities.
@@ A.1 Dataset
As there is no existing dataset specifically for component-controllable personalization, we curate a dataset from the internet to conduct experiments. Unlike previous works (Ruiz et al., 2023; Kumari et al., 2023) that focus on very few categories of concepts, the dataset contains various domains of concepts and components, such as characters, animation, buildings, objects, and animals. Overall, the dataset consists of 23 concept-component pairs totally with 138 reference images, where each concept/component contains 3 reference images and a corresponding category label. The scale of this dataset is aligned with the scale of those datasets used in compared methods (Gal et al., 2022; Ruiz et al., 2023; Kumari et al., 2023; Avrahami et al., 2023; Safaee et al., 2024).
@@ A.2 Implementation
We utilize Stable Diffusion 2.1 (Rombach et al., 2022) as the pretrained T2I diffusion model. As commonly done, the resolution of reference images is set to 512 $\times$ 512. Besides, the rank and alpha of the LoRA module are set to 32. To simplify concept learning, we exclude the region of the target component from the segmentation masks of the target concept, e.g., remove the hair from the person in a “<person> + <hair>” pair. For the warm-up and DS-Bal stage, we set the learning rate to 1e-4 and 1e-5 and the training steps to 200 and 300. Moreover, the learning rate is further scaled by the batch size, which is set to completely contain a concept-component pair. For the cross-attention loss, we follow (Avrahami et al., 2023) to average the corresponding cross-attention maps at resolution 16 $\times$ 16 and normalized them to [0, 1], yielding $A_{\theta}(p_{n},z_{nk}^{(t)})$ in Equation 2. MagicTailor is trained with an AdamW (Loshchilov & Hutter, 2017) optimizer and a DDPM (Ho et al., 2020) sampler on an NVIDIA A100 GPU. For one concept-component pair, it runs for about 5 minutes. All experiments are accomplished with Python 3.10.11 and PyTorch 1.13.1, based on CUDA 11.6. As done in (Avrahami et al., 2023), the tensor precision is set to float16 to accelerate training. For a fair comparison, all random seeds are fixed at 0 in each experiment, and all compared methods use the same implementation above except for method-specific configurations.
A.3 Text Prompts For Evaluation
@@
To generate images for evaluation, we carefully design 20 text prompts covering extensive situations, which are listed in Table 5. These text prompts can be divided into four aspects, including recontextualization, restylization, interaction, and property modification, where each aspect is composed of 5 text prompts. In recontextualization, we change the contexts to different locations and periods. In restylization, we transfer concepts into various artistic styles. In interaction, we explore the spatial interaction with other concepts. In property modification, we modify the properties of concepts in rendering, views, and materials. Such a group of diverse text prompts allows us to systemically evaluate the generalization capability of a method.
A.4 Scheme of Generating Evaluation Images
We generate 32 images per text prompt for each pair, using a DDIM (Song et al., 2020) sampler with 50 steps and a classifier-free guidance scale of 7.5. To ensure fairness, we fix the random seed within the range of [0, 31] across all methods. This process results in a total of 14,720 images for each method to be evaluated, ensuring a robust and thorough comparison.
A.5 Automatic Metrics
We utilize four automatic metrics in the aspects of text alignment (CLIP-T (Gal et al., 2022)) and identity fidelity (CLIP-I (Radford et al., 2021), DINO (Oquab et al., 2023), DreamSim (Fu et al., 2023)). To precisely measure identity fidelity, we improve the traditional measurement approach for personalization. This is because a reference image of the target concept/component could contain an undesired component/concept that is not expected to appear in evaluation images. Specifically, we use Grounded-SAM (Ren et al., 2024) to segment out the concept and component in each reference and evaluation image. Then, we further eliminate the target component from the segmented concept as we have done during training. Such a process is similar to the one adopted in (Avrahami et al., 2023). As a result, using the segmented version of evaluation images and reference images, we can accurately calculate the metrics of identity fidelity.
A.6 User Study
In addition to using automatic metrics, we further evaluate the methods with a user study. Specifically, we design a questionnaire to display 20 groups of evaluation images generated by our method and other methods. Besides, each group also contains the corresponding text prompt and the reference images of the concept and component, where we adopt the same text prompts that are used to calculate CLIP-T. The results of our method and all the compared methods are presented on the same page. Clear rules are established for users to evaluate in three aspects, including text alignment, identity fidelity, and generation quality. Users are requested to select the best result in each group by answering the corresponding questions of these three aspects. We hide all the method names and randomize the order of methods to ensure fairness. Finally, 3,180 valid answers are collected for a sufficient quantitative evaluation of human preferences.
@@ A.6 User Study
A.7 Compared Methods
In our experiments, we compare MagicTailor with SOTA methods in the domain of personalization, including Textual Inversion (TI) (Gal et al., 2022), DreamBooth-LoRA (DB) (Ruiz et al., 2023), Custom Diffusion (CD) (Kumari et al., 2023), Break-A-Scene (BAS) (Avrahami et al., 2023), and CLiC (Safaee et al., 2024). For TI, DB, and CD, we use the third-party implementation in Diffusers [^1] https://huggingface.co/docs/diffusers/index. For BAS, we use the official implementation. For CLiC, we reproduce it following the resource paper as the official code is not released. Unless otherwise specified, method-specific configurations are set up by following their resource papers or Diffusers. We empirically adjust the learning rate of CD and CLiC to 1e-4 and 5e-5 respectively, because they perform very poorly with the original learning rates. For a fair and meaningful comparison, these methods should be adapted to our task setting with minimal modification. Therefore, for those methods adopting a vanilla diffusion loss, we integrate the masked diffusion loss (Equation 1) into them while using the same segmentation masks from MagicTailor.
@@ Appendix B Additional Comparisons
B.1 Comparing with Detailed-Text-Guided Generation
@@
One might be curious about whether component-controllable personalization can be accomplished by providing detailed textual descriptions to the T2I model. To investigate this, we separately feed the reference images of the concept and component into GPT-4o [^2] https://platform.openai.com/docs/models/gpt-4o to obtain detailed textual descriptions for them. The text prompt we used is “Please detailedly describe the <concept/component> of the upload images in a parapraph”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we ask GPT-4o to merge these textual descriptions using natural language, and input them into the Stable Diffusion 2.1 (Rombach et al., 2022) to generate the corresponding images. Some examples for a qualitative comparison are shown in Figure 10. As we can see, such an approach cannot achieve satisfactory results, because it is hard to guarantee that visual semantics can be completely expressed by using the combination of text tokens. In contrast, our MagicTailor is able to accurately learn the desired visual semantics of the concept and component from reference images, and thus lead to consistent and excellent generation in this tough task.
B.2 Comparing with Commercial Models
@@
It is also worth exploring whether existing commercial models, which can understand and somehow generate both text and images by themselves or other integrated tools, are capable of handling component-controllable personalization. We choose two widely recognized commercial models, GPT-4o and Gemini 1.5 Flash [^3] https://gemini.google.com/, for a qualitative comparison. First, we separately feed the reference images of the concept and component into them, along with the text prompt of “The uploaded images contain a special instance of the <concept/component>, please mark it as #<concept/component>”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we instruct them to perform image generation, using the text prompt of “Please generate images containing #<concept> with #<component>”, where “<concept>” and “<component>” are replaced with the category label of the concept and component, respectively. As shown in Figure 11, these models struggle to reproduce the given concept, let alone reconfigure the concept’s component. Whereas, our MagicTailor achieves superior results in component-controllable personalization, using a dedicated framework designed for this task. It demonstrates that, even though large commercial models are able to tackle multiple general tasks, there is also plenty of room for the community to explore specialized tasks for real-world applications.
@@
Appendix C Additional Ablation Studies
C.1 Necessity of Warm-up Training
In MagicTailor, we start with a warm-up phase for the T2I model to preliminarily inject the knowledge for the subsequent phase of DS-Bal. Here we investigate the necessity of such a warm-up phase for generation performance. In Table 6, when removing the warm-up phase, even though MagicTailor could obtain slight improvement in text alignment, it severely suffers from the huge drop in identity fidelity. This is because such a scheme makes it difficult to construct a decent momentum denoising U-Net for DS-Bal. Whereas integrated with a warm-up phase, MagicTailor can achieve superior overall performance, which is attributed to the knowledge reserved from warm-up.
@@ C.1 Necessity of Warm-up Training
C.2 Performance on Different Numbers of Reference Images
As described in Appendix A, each concept/component contains 3 reference images in a concept-component pair of the dataset. Here we change the number of reference images to analyze the performance variation of MagicTailor. The qualitative results are presented in Figure 12. When the number of reference images is reduced, MagicTailor can still show satisfactory performance. This also demonstrates that, while more reference images could lead to better generalization ability, one reference image per concept/component is enough to obtain a decent result with our MagicTailor.
C.3 Robustness on Linking Words
Generally, we use “with” to link the pseudo-words of the concept and component in a text prompt, e.g., “<person> with <beard>, in Von Gogh style”. Here we evaluate the robustness of our method on different linking words. We choose several words, which are commonly used to indicate ownership or association, to construct text prompts and then feed them into the same fine-tuned T2I model. As shown in Figure 13, the generation performance of our MagicTailor remains robust regardless of the linking word used, exhibiting its flexibility to textual descriptions.
Figure 13: Ablation of the linking word. We present qualitative results generated with different linking words, showing the robustness of MagicTailor.
Appendix D More Qualitative Results
In Figure 14, we provide more evaluation images for a substantial qualitative comparison. It can be clearly observed that semantic pollution remains an intractable problem for these compared methods. While employing the masked diffusion loss, they still fall short in suppressing the appearance of undesired visual semantics. This is due to the leak of an effective mechanism to alleviate the T2I model’s perception for these semantics. To address this, our MagicTailor utilizes Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics during the learning phase, and thus achieve better performance. On the other hand, the compared methods are also severely influenced by semantic imbalance, resulting in overemphasis or even overfitting on the concept or component. This is because the inherent imbalance of visual semantics complicates the learning process. In response to this issue, our MagicTailor applies Dual-Stream Balancing (DS-Bal) to balance the learning of visual semantics, effectively showcasing its prowess in this tough task.

### v2

@@ 4.5 Applications
After learning from a concept-component pair, MagicTailor can also enable decoupled generation. As shown in Fig. 7 (a), MagicTailor can generate the target concept and component separately in various and even cross-domain contexts. This should be credited to its remarkable ability to capture different-level visual semantics. Such an ability extends the flexibility of the possible combination between the concept and component.
Controlling Multiple Components.
@@
In this paper, we focus on personalizing one concept and one component, because such a setting is enough to cover extensive scenarios, and can be further extended to reconfigure multiple components with an iterative procedure. However, as shown in Fig. 7 (b), our MagicTailor also exhibits the potential to control two components simultaneously. Handling more components remains a prospective direction of exploring better control over diverse elements for a single concept.
Figure 8: Ablation of loss weights. We report CLIP-T [11] for text alignment, and DreamSim [10] for identity fidelity as it is most similar to human judgments. For reference, we also present the second-best results in Table 1, highlighting our robustness.
Enhancing Other Generative Tools.
@@
In Fig. 7 (c), we provide some interesting method combinations to show that our MagicTailor can enhance other generative tools. The combined tools include ContorlNet [56], CSGO [51], and InstantMesh [52]. As we can see, MagicTailor can be seamlessly integrated into these tools, furnishing them with an additional ability to control the concept’s component in their pipelines. For instance, working with MagicTailor, InstantMesh can conveniently achieve fine-grained design of 3D mesh, demonstrating the practicability of MagicTailor in collaborative applications.
@@ 5 Conclusion
We introduce component-controllable personalization, enabling precise customization of individual components within a concept. Our framework, MagicTailor, uses Dynamic Masked Degradation (DM-Deg) to suppress unwanted semantics and Dual-Stream Balancing (DS-Bal) to ensure balanced learning. Experiments show that MagicTailor sets a new standard in this task, with promising creative applications. In the future, we would like to extend our approach to broader image and video generation, enabling finer control over multi-level visual semantics for enhanced generative capabilities.
Supplementary Material
@@ A.1 Dataset
As there is no existing dataset specifically for component-controllable personalization, we curate a dataset from the internet to conduct experiments. Particularly, unlike previous works [36, 19] that focus on very few categories of concepts, the dataset contains concepts and components from various domains, such as characters, animation, buildings, objects, and animals. Overall, the dataset consists of 23 concept-component pairs totally with 138 reference images, where each concept/component contains 3 reference images and a corresponding category label. The scale of this dataset is aligned with the scale of those datasets used in the compared methods [11, 36, 19, 2, 38].
@@ A.2 Implementation
We utilize Stable Diffusion 2.1 [35] as the pretrained T2I diffusion model. As commonly done, the resolution of reference images is set to 512 $\times$ 512. Besides, the rank and alpha of the LoRA module are set to 32. To simplify concept learning, we exclude the region of the target component from the segmentation masks of the target concept, e.g., remove the hair from the person in a “<person> + <hair>” pair. For the warm-up and DS-Bal stage, we set the learning rate to 1e-4 and 1e-5 and the training steps to 200 and 300. Moreover, the learning rate is further scaled by the batch size, which is set to completely contain a concept-component pair. For the cross-attention loss, we follow [2] to average the corresponding cross-attention maps at resolution 16 $\times$ 16 and normalized them to [0, 1]. The model is trained with an AdamW [23] optimizer and a DDPM [14] sampler. As done in [2], the tensor precision is set to float16 to accelerate training. For a fair comparison, all random seeds are fixed at 0 in each experiment, and all compared methods use the same implementation above except for method-specific configurations.
A.3 Evaluation
@@
To generate images for evaluation, we carefully design 20 text prompts covering extensive situations, which are listed in Tab. 6. These text prompts can be divided into four aspects, including recontextualization, restylization, interaction, and property modification, where each aspect is composed of 5 text prompts. In recontextualization, we change the contexts to different locations and periods. In restylization, we transfer concepts into various artistic styles. In interaction, we explore the spatial interaction with other concepts. In property modification, we modify the properties of concepts in rendering, views, and materials. Such a group of diverse text prompts allows us to systemically evaluate the generalization capability of a method. We generate 32 images per text prompt for each pair, using a DDIM [41] sampler with 50 steps and a classifier-free guidance scale of 7.5. To ensure fairness, we fix the random seed within the range of [0, 31] across all methods. This process results in a total of 14,720 images for each method to be evaluated, ensuring a robust and thorough comparison.
@@
A.4 Automatic Metrics
We utilize four automatic metrics in the aspects of text alignment (CLIP-T [11]) and identity fidelity (CLIP-I [29], DINO [26], DreamSim [10]). To precisely measure identity fidelity, we improve the traditional measurement approach for personalization. This is because a reference image of the target concept/component could contain an undesired component/concept that is not expected to appear in evaluation images. Specifically, we use Grounded-SAM [34] to segment out the concept and component in each reference and evaluation image. Then, we further eliminate the target component from the segmented concept as we have done during training. Such a process is similar to the one adopted in [2]. As a result, using the segmented version of evaluation images and reference images, we can accurately calculate the metrics of identity fidelity.
A.5 User Study
We further evaluate the methods with a user study. Specifically, we design a questionnaire to display 20 groups of evaluation images generated by our method and other methods. Besides, each group also contains the corresponding text prompt and the reference images of the concept and component, where we adopt the same text prompts that are used to calculate CLIP-T. The results of our method and all the compared methods are presented on the same page. Clear rules are established for users to evaluate in three aspects, including text alignment, identity fidelity, and generation quality. Users are requested to select the best result in each group by answering the corresponding questions of these three aspects. We hide all the method names and randomize the order of methods to ensure fairness. Finally, 3,180 valid answers are collected for a sufficient quantitative evaluation of human preferences.
A.6 Compared Methods
In our experiments, we compare MagicTailor with SOTA methods in the domain of personalization, including Textual Inversion (TI) [11], DreamBooth-LoRA (DB) [36], Custom Diffusion (CD) [19], Break-A-Scene (BAS) [2], and CLiC [38]. We select these methods because TI, DB, and CD are three representatives of personalization frameworks and BAS and CLiC are highly relevant to learning fine-grained elements from reference images. For TI, DB, and CD, we use the third-party implementation in Diffusers [8]. For BAS, we use the official implementation. For CLiC, we reproduce it following the resource paper as the official code is not released. Unless otherwise specified, method-specific configurations are set up by following their resource papers or Diffusers. We empirically adjust the learning rate of CD and CLiC to 1e-4 and 5e-5 respectively, because they perform very poorly with the original learning rates. For a fair and meaningful comparison, these methods should be adapted to our task setting with minimal modification. Therefore, for those methods adopting a vanilla diffusion loss, we integrate the masked diffusion loss into them while using the same segmentation masks from MagicTailor.
Figure 10: Compatibility with different backbones. We equip MagicTailor with SD 1.5 [35], SD 2.1 [35], and SDXL [27]. The results show that MagicTailor can be generalized to multiple backbones, and a better backbone could provide better performance.
@@ Appendix B Additional Comparisons
B.1 Detailed Text-Guided Generation
@@
One might wonder if component-controllable personalization can be accomplished by providing detailed textual descriptions to the T2I model. To investigate this, we separately feed the reference images of the concept and component into GPT-4o [17] to obtain detailed textual descriptions for them. The text prompt we used is “Please detailedly describe the <concept/component> of the upload images in a parapraph”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we ask GPT-4o to merge these textual descriptions using natural language, and input them into the Stable Diffusion 2.1 [35] to generate the corresponding images. Some examples for a qualitative comparison are shown in Fig. 9. As we can see, such an approach cannot achieve satisfactory results, because it is hard to guarantee that visual semantics can be completely expressed by using the combination of text tokens. In contrast, our MagicTailor is able to accurately learn the desired visual semantics of the concept and component from reference images, and thus lead to consistent and excellent generation in this tough task.
B.2 Commercial Models
@@
It is also worth exploring whether existing commercial models, which can understand and somehow generate both text and images by themselves or other integrated tools, are capable of handling component-controllable personalization. We choose two widely recognized commercial models, GPT-4o [17] and Gemini 1.5 Flash [43], for a qualitative comparison. First, we separately feed the reference images of the concept and component into them, along with the text prompt of “The uploaded images contain a special instance of the <concept/component>, please mark it as #<concept/component>”, where “<concept/component>” is replaced with the category label of the concept or component. Then, we instruct them to perform image generation, using the text prompt of “Please generate images containing #<concept> with #<component>”, where “<concept>” and “<component>” are replaced with the category label of the concept and component, respectively. As shown in Fig. 12, these models struggle to reproduce the given concept, let alone reconfigure the concept’s component. Whereas, our MagicTailor achieves superior results in component-controllable personalization, using a dedicated framework designed for this task. It demonstrates that, even though large commercial models are able to tackle multiple general tasks, there is also plenty of room for the community to explore specialized tasks for real-world applications.
Figure 11: Effectiveness on difficult training pairs. We show the results of two hard cases involving large geometric discrepancy and cross-domain interactions, showing that MagicTailor can effectively handle such challenging scenarios.
Appendix C Generalizability Analysis
As discussed in Sec. A, we follow a strict and standardized setup to conduct performance evaluation for all the methods in our experiments. In this section, we perform a comprehensive generalizability analysis for MagicTailor by extending the setup to more flexible scenarios. Specifically, the analysis focuses on the model perspective, where we explore MagicTailor’s compatibility with different backbones (Sec. C.1), and the data perspective, where we study its effectiveness on difficult training pairs (Sec. C.2) and complex text prompts (Sec. C.3).
@@ Appendix C Generalizability Analysis
Figure 13: Effectiveness on complex text prompts. We present qualitative results generated with complex text prompts. In addition to those well-categorized text prompts, our MagicTailor can also follow more complex ones to generate text-aligned images.
C.1 Compatibility with Different Backbones
We use Stable Diffusion 2.1 (SD 2.1) [35], which is widely adopted by the compared methods, as our backbone in the main paper. It is worth noting that our MagicTailor is a general method also suitable for other T2I diffusion models, because only input images and the denoising paradigm are reformulated. To evaluate the compatibility, we employ MagicTailor in other backbones, including Stable Diffusion 1.5 (SD 1.5) [35] and Stable Diffusion XL (SDXL) [27]. As shown in Fig. 10, our MagicTailor can also achieve satisfactory results when cooperating with various backbones. Notably, we only change the LoRA rank for adapting to different learning capabilities, and directly use the original hyperparameter values, which shows the remarkable generalizability of MagicTailor.
C.2 Effectiveness on Difficult Training Pairs
Appropriate processing for reference images, such as filtering, scaling, and cropping, can help to construct a better training pair for personalization. Unfortunately, not all training pairs are ideal in real-world scenarios. Therefore, here we investigate the performance of MagicTailor when handling difficult training pairs. Specifically, we consider two hard cases. The first is a pair with large geometric discrepancy, such as “<person>” in an upper body portrait and “<hair>” in a profile photo. The second is a pair with cross-domain interactions, such as “<person>” and “<ear>” of dogs. As shown in Fig. 11, even facing these hard cases, MagicTailor is also capable of personalizing the target concept and component with high fidelity.
C.3 Effectiveness on Complex Text Prompts
In our experiments, we use well-categorized text prompts in four aspects, including recontextualization, restylization, interaction, and property modification, for qualitative and quantitative comparisons. Here we further evaluate MagicTailor’s performance on other complex text prompts involving more complicated contexts and elements. As shown in Fig. 13, our MagicTailor can also effectively generate text-aligned images when performing fidelity personalization, showing its ability to handle diverse user needs.
Figure 14: Ablation of DM-Deg via replacement with SID. We compare our DM-Deg with SID [18] that aims to produce informative prompts for training. Besides, we also present baseline (i.e., removing DM-Deg from MagicTailor) results for reference. This comparison indicates the effectiveness of our DM-Deg in addressing semantic pollution.
Appendix D Additional Ablation Studies
D.1 Effectiveness of DM-Deg over Using Informative Training Prompts
One might be curious about whether it is not necessary to employ the proposed DM-Deg, but perhaps to use informative text prompts during training to provide textual prior knowledge for learning the target concept and component. To investigate this, we use Selectively Informative Description (SID) [18] with GPT-4o [17] to construct text prompts for the target concept and component, and then use them for training. As shown in Fig. 14, such an approach cannot address semantic pollution well, where unwanted visual semantics still affect the personalized concept. In contrast, DM-Deg effectively prevents semantic pollution by dynamically perturbing those undesired visual semantics, verifying its remarkable significance in this task.
D.2 Necessity of Warm-Up Training
In MagicTailor, we start with a warm-up phase for the T2I model to preliminarily inject the knowledge for the subsequent phase of DS-Bal. Here we investigate the necessity of such a warm-up phase for generation performance. In Tab. 7, when removing the warm-up phase, even though MagicTailor could obtain slight improvement in text alignment, it severely suffers from the huge drop in identity fidelity. This is because such a scheme makes it difficult to construct a decent momentum denoising U-Net for DS-Bal. Whereas integrated with a warm-up phase, MagicTailor can achieve superior overall performance, which is attributed to the knowledge reserved from warm-up.
@@ D.2 Necessity of Warm-Up Training
D.3 Performance on Different Numbers of Reference Images
As described in Appendix A, each concept/component contains 3 reference images in a concept-component pair of the dataset. Here we change the number of reference images to analyze the performance variation of MagicTailor. The qualitative results are presented in Fig. 15. When the number of reference images is reduced, MagicTailor can still show satisfactory performance. This also demonstrates that, while more reference images could lead to better generalization ability, one reference image per concept/component is enough to obtain a decent result with our MagicTailor.
D.4 Robustness on Linking Words
Generally, we use “with” to link the pseudo-words of the concept and component in a text prompt, e.g., “<person> with <beard>, in Von Gogh style”. Here we evaluate the robustness of our method on different linking words. We choose several words, which are commonly used to indicate ownership or association, to construct text prompts and then feed them into the same fine-tuned T2I model. As shown in Fig. 16, the generation performance of our MagicTailor remains robust regardless of the linking word used, exhibiting its flexibility to textual descriptions.
Figure 16: Ablation of linking words. We present qualitative results generated with different linking words, demonstrating the robustness of MagicTailor.
Appendix E More Qualitative Results
In Fig. 17 & Fig. 18, we provide more evaluation images for a substantial qualitative comparison. It can be clearly observed that semantic pollution remains an intractable problem for these compared methods. This is due to the leak of an effective mechanism to alleviate the T2I model’s perception for these semantics. To address this, our MagicTailor utilizes DM-Deg to dynamically perturb undesired visual semantics during the learning phase, and thus achieve better performance. On the other hand, the compared methods are also severely influenced by semantic imbalance, resulting in overemphasis or even overfitting on the concept or component. This is because the inherent imbalance of visual semantics complicates the learning process. In response to this issue, our MagicTailor applies DS-Bal to balance the learning of visual semantics, effectively showcasing its prowess in this tough task.